# Real-time 3D-aware Portrait Editing from a Single Image

### v2

@@ Abstract
This work presents 3DPE, a practical method that can efficiently edit a face image following given prompts, like reference images or text descriptions, in a 3D-aware manner. To this end, a lightweight module is distilled from a 3D portrait generator and a text-to-image model, which provide prior knowledge of face geometry and superior editing capability, respectively. Such a design brings two compelling advantages over existing approaches. First, our system achieves real-time editing with a feedforward network (i.e., $\sim$0.04s per image), over 100$\times$ faster than the second competitor. Second, thanks to the powerful priors, our module could focus on the learning of editing-related variations, such that it manages to handle various types of editing simultaneously in the training phase and further supports fast adaptation to user-specified customized types of editing during inference (e.g., with $\sim$5min fine-tuning per style). The code, the model, and the interface will be made publicly available to facilitate future research.
@@ 1 Introduction
Inferring the geometry and appearance from a single-view portrait image has become mature and practical [42, 37, 74, 65, 72, 68, 22], largely attributed to the utilization of priors in various 2D/3D generative models. However, only performing geometry reconstruction is insufficient. The significance of 3D editing of portraits, driven by user intentions, and the need for streamlined efficiency in the editing process has been steadily increasing. This is particularly crucial in real-world applications such as AR/VR, 3D telepresence, and video conferencing, where real-time editing is often essential. Consequently, a key question arises: How can we effectively address the challenge of attaining high-fidelity portrait editing while ensuring real-time efficiency?
@@
To this end, we present 3DPE, a real-time 3D-aware Portrait Editing method driven by user-defined prompts. As shown in Fig. 1, when provided with a single-view portrait image, our approach empowers a versatile range of editing styles with one model through flexible instructions, including images and texts. We leverage the powerful 3D prior from a 3D-aware face generator and achieve a high-fidelity 3D reconstruction of the portrait image. Subsequently, we distill editing knowledge from a text-guided image editing model into a lightweight module integrated with the 3D-aware generator. This module, characterized by its minimal computational cost, allows our system to maintain real-time inference and excel in handling various types of editing, all the while ensuring good 3D consistency. An additional advantage is that our model supports customization through user-specified prompts with fast adaptation speed. This empowers users to build their own editing model at a minimal cost, enabling our system to cater to a broader audience.
Our system achieves real-time 3D-aware portrait editing through the utilization of a feedforward network, with a processing time of 40ms on a standard consumer GPU. Additionally, we present a comprehensive evaluation of our method using various prompts both quantitatively and qualitatively. Our design choices are also validated through comparisons with ablated variants of our method. Compared with baseline methods, our approach demonstrates superior 3D consistency, precise texture alignment, and a substantial improvement in inference time, as reflected by the evaluation metrics. We demonstrate the versatility of our method by showcasing its capacity to perform a wide variety of edits, including text and image prompts, on portrait images. The code, the model, and the interface will be made publicly available.
@@ 2.1 Generative Face Priors
Generative models aim at modeling the underlying distribution of the data, containing a wealth of prior knowledge. Recently, 3D GANs [57, 14, 13, 75, 64, 58, 63, 62, 48, 25, 46, 69] are mostly adopted to learn 3D faces from single-view image dataset. The interior rich domain-specific geometry priors enable various face-related applications such as image editing [68, 9, 66, 65, 83, 77] and domain adaptation [34, 36, 80, 1]. Large-scale diffusion models [55], in contrast, encode knowledge of huge datasets and thus can provide general prior information. Such priors are broadly leveraged for tasks such as image editing [45, 81, 24, 7, 10], customization [56, 31, 44, 38], and video editing [47, 43, 11, 52, 71, 78, 23, 12]. In our approach, we capitalize on the advantages of both the geometry prior derived from 3D GANs and the broader editing prior offered by large-scale text-to-image diffusion models, instead of relying solely on a single type of prior.
@@ 3 Method
Our approach takes a portrait image $\mathbf{I}$ and its camera pose c, obtained via the face pose estimator [18]. Additionally, the referenced prompts $\mathbf{P}$ (images or texts) serve as the editing instructions. The outcome of our model is an edited version of the 3D portraits characterized by NeRF in accordance with the prompts. The edited image is denoted as $\mathbf{I}_{p}$. Within our framework, we achieve this objective by leveraging the knowledge of 3D GANs and a text-guided image editing model, which is distilled into a lightweight module. Following distillation, our method enables real-time 3D portrait editing and efficient adaptation to user-specified prompts.
In Sec. 3.1, we offer an illustration of 3D GANs and diffusion models. The process of distilling these priors into a lightweight module is outlined in Sec. 3.2. Finally, the details for model training and inference, along with the fast adaptation for customized prompts, are presented in Sec. 3.3.
@@ 3.1 Preliminary
3D GAN Prior for Portrait Reconstruction. The 3D-aware GANs showcase the ability to synthesize photorealistic 3D images using a collection of single-view images. Notably, EG3D [13] introduces an efficient triplane 3D representation, demonstrating high-quality 3D-aware image rendering. Once trained, the generator of EG3D can be applied for single-image 3D reconstruction via GAN inversion. However, existing 3D GAN inversion methods often encounter geometry distortion or exhibit slow inference speed. To tackle these challenges, we utilize Live3D [68], a state-of-the-art single-image portrait 3D reconstruction model built upon EG3D, preserving geometry quality while ensuring real-time performance. Live3D employs a two-branch encoder, $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$, to extract different resolution features from the input image. It then utilizes a ViT-based $\mathrm{E}_{t}(\cdot)$ decoder [19] to transform the fused encoder features into the triplane $\mathbf{T}$:
@@
This triplane $\mathbf{T}$ is subsequently used in conjunction with the volume rendering module and upsampler of EG3D to generate photorealistic view synthesis given the camera pose $\mathrm{c}$. For clarification, we use $\mathrm{R}(\cdot)$ to denote this process.
Diffusion Prior for Portrait Editing. Only performing 3D reconstruction is insufficient as our goal is to edit portrait images. Large-scale diffusion models [55], trained on vast text-image pairs, can synthesize realistic photos with text input, offering powerful editing prior. Despite their effectiveness for face images, the editing process is often slow due to step-by-step optimization or multiple inferences in diffusion models. Moreover, obtaining a large amount of high-quality paired data with multi-view consistency from the diffusion model is extremely challenging, making the application of 3D-aware editing from a single image impractical. Therefore, our goal is to distill editing knowledge from the diffusion model, integrate it with the 3D prior in Live3D into a lightweight module, and employ it for real-time portrait editing.
@@ 4.1 Experimental Setup
Training Settings. For real faces, we adopt the FFHQ dataset [35] at $512\times 512$ resolution with camera parameters aligned by EG3D [13]. In order to obtain the stylized images as pseudo labels, and the corresponding textual and visual prompts, we leverage the text-guided editing model InstructPix2Pix (IP2P) [7] to edit the real faces. Specifically, we manually generate and select textual prompts, and send them to IP2P together with the real images to obtain the edited 2D images. The text prompts of our training and testing set are obtained by augmenting the aforementioned textual prompts with the large language model GPT [8] following IP2P [7]. While the image prompts of our training and testing set are picked from the edited 2D images. The dataset details and samples are provided in the appendix for further investigation. We conduct experiments on 20 styles, and for each style, we synthesize 1000 images, resulting in 20,000 image pairs with text prompts for model training. For each style, we use 8 textual prompts for training and 5 textual prompts for testing. We adopt a learning rate of 5e-5 and optimize the model for 60k iterations with a batch size of 32. The entire training procedure is completed over a period of 40 hours utilizing 8 NVIDIA A100 GPUs.
@@
Baselines. We conduct comparisons against two categories of methods that achieve analogous outcomes: 1) 3D reconstruction coupled with 3D editing; 2) 2D editing followed by 3D reconstruction. For the first category, we implement three baseline approaches, wherein we firstly leverage PTI [54] or Live3D [68] for reconstruction and edit both of them following StyleCLIP [49]. Specifically, for experiments in PTI + StyleCLIP, we first conduct pivotal tuning [54] based on the predicted latent codes predicted by e4e [67] to reconstruct the given image in a 3D-aware manner. Then we perform editing on the latent codes following [49], where CLIP loss [53] between the model output and the prompt (no matter it is an image or text) is utilized. The setting of Live3D + StyleCLIP is similar to PTI + StyleCLIP except that we obtain the triplane by Live3D and perform editing on this 3D representation. SPI [79] + StyleGAN-NADA [21] (StyleNADA for abbreviation in Fig. 4) serves as the third baseline in the first category, where SPI is used for 3D reconstruction and StyleGAN-NADA is utilized to optimize the 3D generator for editing. In the second category, we employ a diffusion-based text-guided 2D image editing method IP2P [7] prior to executing 3D reconstruction via Live3D. IP2P being a text-instructed editing model, we need to first extract text prompts from the reference image with the image caption algorithm BLIP [41] when image prompts are input. The 2D edited results then could be obtained by sending the input image and the text prompt to IP2P. The final 3D results are acquired by inverting the 2D editing results with Live3D.
@@
Figure 5: The learning process and results of customized prompt adaptation. For the novel style golden statue instructed by reference images, we show the intermediate testing results at 10s, 1min, 2min, and 5min during adaptation with 10 paired training images.
@@ 4.2 Efficient 3D-aware Portrait Editing
We make an in-depth analysis of our efficient portrait editing system both quantitatively and qualitatively. The results are included in Tab. 1 and Fig. 4. For reference alignment, we do not report $CLIP_{r}$ for methods that leverage CLIP for optimization since it is evaluated with CLIP as well. The standout feature of our system is its efficiency. Our method achieves an inference speed of merely 40ms, which improves over 100 times compared to the fastest existing baselines, which require around 10 seconds. Because of the efficient knowledge distillation, our approach is also good in preserving the identity, adhering more closely to the reference prompts, and achieving the best 3D consistency.
Compared with PTI + CLIP methods, our model can enable precise alignment with input and prompts, while the baseline always generates low-fidelity textures, and the resulting geometry has many artifacts. Although the Live3D + CLIP achieves the best $ID_{t}$ score, the edited results are basically unchanged compared to the input. We suspect that Live3D does not have a latent space and loses the editing priors, making it challenging for the results of CLIP optimization to align well with prompts. SPI + StyleGAN-NADA performs relatively slowly (18 minutes per instance) because both the 3D reconstruction and the editing are achieved by optimizing the generator. Although it can produce realistic textures, it fails to reconstruct fine details and to edit following the prompts. In contrast to the 2D editing and subsequent 3D reconstruction pipeline (IP2P + Live3D), our system produces textures consistent with the prompt and aligns better with the input structure. As demonstrated by the instance in the first column, IP2P + Live3D introduces texture artifacts (the secondary face) due to the unstable editing performance of IP2P, while our distillation-based method excels in this aspect. As the second sample in Fig. 4 shows, our model can consistently retain the hair and cloth structure, while IP2P + Live3D cannot handle this. The image caption result by BLIP of the third case turns out to be “a painting of a woman with blue eyes and a red headband”. This sample showcases that, the method of IP2P + Live3D could not handle well editing guided by image prompts due to the improper image caption sent to the 2D editing model, especially for the styles hard to describe with texts. The caption would integrate the content and the desired style of the reference image, which deteriorates both the editing and the identity preservation.
@@ 4.3 Adaptation for Customized Editing
With the trained editing network, our system already supports a variety of 3D-aware editing styles. Following DreamBooth [56] and to expand the selection and better conform to users’ preferences, we offer an efficient method for fast adaptation to the user-defined customized prompts. Users can personalize the editing network by providing a modest set of 10 reference editing pairs. These pairs can be either handpicked from artist-created examples or produced using text-guided image editing models. As described in Sec. 3.3, only the partial of the encoder would be optimized. The adaptation process itself is remarkably fast, requiring only about 5 minutes to accomplish the learning of the customized prompts. Please also note that, once the user-defined style has been trained, all the incoming testing results can be obtained in a real-time manner. In Fig. 5, we present the intermediate testing results in the adaptation process. Our model can quickly master the customized knowledge in about 2 minutes. With further training (e.g., 5 minutes), the edited results become more stylized, demonstrating a trade-off between authenticity and stylization. Upon completion of the adaptation, the system allows users to edit testing inputs in these newly learned styles with a minimal inference time of 0.04s. This rapid performance indicates that the system is well-optimized for real-time applications, providing a seamless and efficient user experience.
@@ 4.4 Ablation Studies
Ablation on Tuning Data Scale for Customized Prompt Adaptations. To investigate the ability of our method to adapt to user-defined novel prompts, we conduct an ablation study on the number of novel-prompt data pairs required for training. Similar to the prior experiments on the pretraining scale, we perform adaptation with 2, 5, 10, 20, or 50 pairs of novel-prompt data on an A6000 GPU. We experimentally find that our model can converge in about 5 minutes. Therefore, we report the perceptual distance between the model output and the provided ground truth at 0.1min, 1min, 2min, and 5min for comparison. As in Tab. 3, our model can adapt to the new style in around 2 minutes. As the number of provided data pairs increases, the model can adapt faster at the beginning.
@@ 4.5 Applications
One application of our pipeline is to edit videos, requiring accurate reconstruction and efficient editing simultaneously. We demonstrate the effectiveness of our pipeline on talking face videos, as illustrated in Fig. 7. In this experiment, we perform sequence editing in a frame-by-frame manner by sending an image sequence from TalkingHead-1KH [70] dataset together with the text prompt "Make him a bronze statue" to 3DPE. The three rows in Fig. 7 respectively illustrate the input frames, the edited frames, and the novel views. Our method accurately reconstructs facial expressions, while achieving high-quality editing results. Since our method is 3D-aware, we can naturally view videos from novel perspectives. To better experience our method, we design an interactive system that allows users for interactive editing, which is shown in the appendix. Users are required to provide the image or text prompt to indicate the desired style and input the image that is to be edited, and the system will apply our method to output the edited image as well as novel-view edited results.

### v3

@@ Abstract
This work presents 3DPE, a practical method that can efficiently edit a face image following given prompts, like reference images or text descriptions, in a 3D-aware manner. To this end, a lightweight module is distilled from a 3D portrait generator and a text-to-image model, which provide prior knowledge of face geometry and superior editing capability, respectively. Such a design brings two compelling advantages over existing approaches. First, our method achieves real-time editing with a feedforward network (i.e., $\sim$0.04s per image), over 100$\times$ faster than the second competitor. Second, thanks to the powerful priors, our module could focus on the learning of editing-related variations, such that it manages to handle various types of editing simultaneously in the training phase and further supports fast adaptation to user-specified customized types of editing during inference (e.g., with $\sim$5min fine-tuning per style). Project page can be found here.
@@ 1 Introduction
Inferring the geometry and appearance from a single-view portrait image has become mature and practical [42, 37, 76, 66, 74, 70, 22], largely attributed to the utilization of priors in various 2D/3D generative models. However, only performing geometry reconstruction is insufficient. The significance of 3D portrait editing, driven by user intentions, and the need for streamlined efficiency in the editing process has been steadily increasing. This is particularly crucial in real-world applications such as AR/VR, 3D telepresence, and video conferencing, where real-time editing is often essential. Consequently, a key question arises: How can we effectively address the challenge of attaining high-fidelity portrait editing while ensuring real-time efficiency?
@@
To this end, we present 3DPE, a real-time 3D-aware Portrait Editing method driven by user-defined prompts. As shown in Fig. 1, when provided with a single-view portrait image, our approach empowers a versatile range of editing styles with one model through flexible instructions, including images and texts. We leverage the powerful 3D prior from a 3D-aware face generator and achieve a high-fidelity 3D reconstruction of the portrait image. Subsequently, we distill editing knowledge from a text-guided image editing model into a lightweight module integrated with the 3D-aware generator. This module, characterized by its minimal computational cost, allows our method to maintain real-time inference and excel in handling various types of editing, while ensuring good 3D consistency. An additional advantage is that our model supports customization through user-specified prompts with fast adaptation speed. This empowers users to build their own editing model at a minimal cost, enabling our method to cater to a broader audience.
Our method achieves real-time 3D-aware portrait editing through the utilization of a feedforward network, with a processing time of 40ms on a standard consumer GPU. Additionally, we present a comprehensive evaluation of our method using various prompts both quantitatively and qualitatively. Our design choices are also validated through comparisons with ablated variants of our method. Compared with baseline methods, our approach demonstrates superior 3D consistency, precise texture alignment, and a substantial improvement in inference time, as reflected by the evaluation metrics. We demonstrate the versatility of our method by showcasing its capacity to perform a wide variety of edits, including text and image prompts, on portrait images.
@@ 2.1 Generative Face Priors
Generative models aim at modeling the underlying distribution of the data, containing a wealth of prior knowledge. Recently, 3D GANs [58, 14, 13, 77, 65, 59, 64, 63, 49, 25, 47, 71, 69] are mostly adopted to learn 3D faces from single-view image dataset. The interior rich domain-specific geometry priors enable various face-related applications such as image editing [70, 9, 67, 66, 85, 79] and domain adaptation [34, 36, 82, 1, 46]. Large-scale diffusion models [56], in contrast, encode knowledge of huge datasets and thus can provide general prior information. Such priors are broadly leveraged for tasks such as image editing [45, 83, 24, 7, 10], customization [57, 31, 44, 38], and video editing [48, 43, 11, 53, 73, 80, 23, 12]. In our approach, we capitalize on the advantages of both the geometry prior derived from 3D GANs and the broader editing prior offered by large-scale text-to-image diffusion models, instead of relying solely on a single type of prior.
@@ 3 Method
Our approach takes a portrait image $\mathbf{I}$ and its camera pose c, obtained via the face pose estimator [18]. Additionally, the referenced prompts $\mathbf{P}$ (images or texts) serve as the editing instructions. The outcome of our model is an edited version of the 3D portraits characterized by NeRF in accordance with the prompts. The edited image is denoted as $\mathbf{I}_{p}$. Within our framework, we achieve this objective by leveraging the knowledge of 3D GANs and a text-guided image editing model, which is distilled into a lightweight module. Following distillation, our method enables real-time 3D portrait editing and efficient adaptation to user-specified prompts. In Sec. 3.1, we offer an illustration of 3D GANs and diffusion models. The process of distilling these priors into a lightweight module is outlined in Sec. 3.2. Finally, the details for model training and inference, along with the fast adaptation for customized prompts, are presented in Sec. 3.3.
@@ 3.1 Preliminary
3D GAN Prior for Portrait Reconstruction. The 3D-aware GANs showcase the ability to synthesize photorealistic 3D images using a collection of single-view images. Notably, EG3D [13] introduces an efficient triplane 3D representation, demonstrating high-quality 3D-aware image rendering. Once trained, the generator of EG3D can be applied for single-image 3D reconstruction via GAN inversion. However, existing 3D GAN inversion methods often encounter geometry distortion or exhibit slow inference speed. To tackle these challenges, we utilize Live3D [70], a state-of-the-art single-image portrait 3D reconstruction model built upon EG3D, preserving geometry quality while ensuring real-time performance. Live3D employs a two-branch encoder, $\mathrm{E}_{high}(\cdot)$ and $\mathrm{E}_{low}(\cdot)$, to extract different resolution features from the input. It then utilizes a ViT-based $\mathrm{E}_{t}(\cdot)$ decoder [19] to transform the fused encoder features into the triplane $\mathbf{T}$:
@@
This triplane $\mathbf{T}$ is subsequently used in conjunction with the volume rendering module and upsampler of EG3D to generate photorealistic view synthesis given the camera pose $\mathrm{c}$. For clarity, we use $\mathrm{R}(\cdot)$ to denote this process.
Diffusion Prior for Portrait Editing. Only performing 3D reconstruction is insufficient as our goal is to edit portrait images. Large-scale diffusion models [56], trained on vast text-image pairs, can synthesize realistic photos with text input, offering powerful editing prior. Despite their effectiveness for face images, the editing process is often slow due to step-by-step optimization or multiple inferences in diffusion models. The scarcity of high-quality and multi-view consistent paired data significantly hinders the practical application of 3D-aware editing from a single image, which is hard to acquire with diffusion models. Therefore, our goal is to distill editing knowledge from the diffusion model, integrate it with the 3D prior in Live3D into a lightweight module, and employ it for real-time portrait editing.
@@ 4.1 Experimental Setup
Training Settings. For real faces, we adopt the FFHQ dataset [35] at $512\times 512$ resolution with camera parameters aligned by EG3D [13]. In order to obtain the stylized images as pseudo labels, and the corresponding textual and visual prompts, we leverage the text-guided editing model InstructPix2Pix (IP2P) [7] to edit the real faces. Specifically, we manually generate and select textual prompts, and send them to IP2P together with the real images to obtain the edited 2D images. The text prompts of our training and testing set are obtained by augmenting the aforementioned textual prompts with the large language model GPT [8] following IP2P [7]. While the image prompts of our training and testing set are picked from the edited 2D images. We conduct experiments on 20 styles, and for each style, we synthesize 1000 images, resulting in 20,000 image pairs with text prompts for model training. For each style, we use 8 textual prompts for training and 5 textual prompts for testing. We adopt a learning rate of 5e-5 and optimize the model for 60k iterations with a batch size of 32. The entire training procedure is completed over a period of 40 hours using 8 NVIDIA A100 GPUs.
@@
Baselines. We conduct comparisons against two categories of methods that achieve analogous outcomes: 1) 3D reconstruction coupled with 3D editing; 2) 2D editing followed by 3D reconstruction. For the first category, we implement three baseline approaches, wherein we firstly leverage PTI [55] or Live3D [70] for reconstruction and edit both of them following StyleCLIP [50]. Specifically, for experiments in PTI + StyleCLIP, we first conduct pivotal tuning [55] based on the predicted latent codes predicted by e4e [68] to reconstruct the given image in a 3D-aware manner. Then we perform editing on the latent codes following [50], where CLIP loss [54] is utilized between the model output and the prompt (no matter whether it is an image or text). The setting of Live3D + StyleCLIP is similar to PTI + StyleCLIP except that we obtain the triplane by Live3D and perform editing on this 3D representation. SPI [81] + StyleGAN-NADA [21] (StyleNADA for the abbreviation in Fig. 4) serves as the third baseline in the first category, where SPI is used for 3D reconstruction and StyleGAN-NADA is utilized to optimize the 3D generator for editing. In the second category, we employ a diffusion-based text-guided 2D image editing method IP2P [7] prior to executing 3D reconstruction via Live3D. IP2P being a text-instructed editing model, we need to first extract text prompts from the reference image with the image caption algorithm BLIP [41] when image prompts are input. The 2D edited results could then be obtained by sending the input image and the text prompt to IP2P. The final 3D results are acquired by inverting the 2D editing results with Live3D.
@@
Figure 5: (a) Generated results of customized prompt adaptation and (b) its learning process. We show the intermediate testing results at 10s, 1min, 2min and 5min during adaptation for the novel style golden statue.
@@ 4.2 Efficient 3D-aware Portrait Editing
We make an in-depth analysis of our efficient portrait editing method both quantitatively and qualitatively. The results are included in Tab. 1 and Fig. 4. For reference alignment, we do not report $CLIP_{r}$ for methods that leverage CLIP for optimization since it is evaluated with CLIP as well. The standout feature of our method is its efficiency. Our method achieves an inference speed of merely 40ms, which improves over 100 times compared to the fastest existing baselines, which require around 10 seconds. Because of the efficient knowledge distillation, our approach is also good in preserving the identity, adhering more closely to the reference prompts and achieving the best 3D consistency.
As in Fig. 4, compared with PTI + CLIP methods, our model enables precise alignment with inputs and prompts, while the baseline always generates low-fidelity textures, and the resulting geometry has many artifacts. Although Live3D + CLIP achieves the best $ID_{t}$ score, the edited results are basically unchanged compared to the input. We suspect that Live3D does not possess a latent space and loses the editing priors, making it challenging for the results of CLIP optimization to align well with prompts. SPI + StyleGAN-NADA performs relatively slowly (18 minutes per instance) because both the 3D reconstruction and the editing are achieved by optimizing the generator. Although it can produce realistic textures, it fails to reconstruct fine details and to edit following the prompts. In contrast to the 2D editing and subsequent 3D reconstruction pipeline (IP2P + Live3D), our method produces textures more consistent with the prompt and aligns better with the input structure.
@@ 4.3 Adaptation for Customized Editing
With the trained editing network, our method already supports a variety of 3D-aware editing styles. Following DreamBooth [57] and to expand the selection and better conform to users’ preferences, we offer an efficient method for fast adaptation to the user-defined customized prompts. Users can personalize the editing network by providing a modest set of 10 reference editing pairs. These pairs can be either handpicked from artist-created examples or produced using text-guided image editing models. As described in Sec. 3.3, only the partial of the encoder would be optimized. The adaptation process itself is remarkably fast, requiring only about 5 minutes to accomplish the learning of the customized prompts. Please also note that, once the user-defined style has been trained, all the incoming testing results can be obtained in a real-time manner. In Fig. 5, we present the testing results and learning process of the adaptation. Our model can quickly master the customized knowledge in about 2 minutes. With further training (e.g., 5 minutes), the edited results become more stylized, demonstrating a trade-off between authenticity and stylization. Upon completion of the adaptation, the method allows users to edit testing inputs in these newly learned styles with a minimal inference time of 0.04s. This rapid performance indicates that the method is well-optimized for real-time applications, providing a seamless and efficient user experience.
@@ 4.4 Ablation Studies
Ablation on Tuning Data Scale for Customized Prompt Adaptations. To investigate the ability of our method to adapt to user-defined novel prompts, we conduct an ablation study on the number of novel-prompt data pairs required for training. Similar to the prior experiments on the pretraining scale, we perform adaptation with 2, 5, 10, 20, or 50 pairs of novel-prompt data on a single A6000 GPU. We experimentally find that our model can converge in about 5 minutes, and therefore, we report the perceptual distance between the model output and the provided ground truth at 0.1min, 1min, 2min, and 5min for comparison. As shown in Tab. 3, our model can adapt to the new style in around 2 minutes. As the number of provided data pairs increases, the model can adapt faster at the beginning.
@@ 4.5 Applications
One application of our pipeline is to edit videos, requiring accurate reconstruction and efficient editing simultaneously. We demonstrate the effectiveness of our pipeline on talking face videos, as illustrated in Fig. 7. In this experiment, we perform sequence editing in a frame-by-frame manner by sending an image sequence from TalkingHead-1KH [72] dataset together with the text prompt "Make him a bronze statue" to 3DPE. The three rows in Fig. 7 respectively illustrate the input frames, the edited frames, and the novel views. Our method accurately reconstructs facial expressions, while achieving high-quality editing results. Since our method is 3D-aware, we can naturally view videos from novel perspectives. To better experience our method, we design an interactive system that allows users for interactive editing. Users are required to provide the image or text prompt to indicate the desired style and input an input image that is to be edited, and the system will apply our method on the given image and prompt to output the edited image as well as novel-view edited results.